# Residual Off-Policy RL for Finetuning Behavior Cloning Policies

### v1

@@ Abstract
Recent advances in behavior cloning (BC) have enabled impressive visuomotor control policies. However, these approaches are limited by the quality of human demonstrations, the manual effort required for data collection, and the diminishing returns from increasing offline data. In comparison, reinforcement learning (RL) trains an agent through autonomous interaction with the environment and has shown remarkable success in various domains. Still, training RL policies directly on real-world robots remains challenging due to sample inefficiency, safety concerns, and the difficulty of learning from sparse rewards for long-horizon tasks, especially for high-degree-of-freedom (DoF) systems. We present a recipe that combines the benefits of BC and RL through a residual learning framework. Our approach leverages BC policies as black-box bases and learns lightweight per-step residual corrections via sample-efficient off-policy RL. We demonstrate that our method requires only sparse binary reward signals and can effectively improve manipulation policies on high-degree-of-freedom (DoF) systems in both simulation and the real world. In particular, we demonstrate, to the best of our knowledge, the first successful real-world RL training on a humanoid robot with dexterous hands. Our results demonstrate state-of-the-art performance in various vision-based tasks, pointing towards a practical pathway for deploying RL in the real world.
@@ II RELATED WORK
Behavior Cloning. BC has shown success in various settings, from autoregressive next-token prediction [1, 24, 37, 38] and diffusion policies [4, 39, 6, 40, 7], to large transformers that directly output robot actions [5, 41, 42]. Recent advances [43, 8, 9] leverage large transformers, diffusion and flow matching heads, and action chunking to handle long-horizon tasks and achieve impressive performance. Despite these advances, BC has fundamental limitations. The manual effort and infrastructure needed to scale up data collection to these levels are extremely high, and recent work has shown diminishing returns and performance plateaus with increasing data [6, 13, 12].
Fine-tuning BC Policies with RL. RL fine-tuning has been explored as a way to improve BC policies by letting the agent interact directly with the environment. However, directly fine-tuning modern BC architectures with RL presents significant challenges. State-of-the-art BC policies typically use action-chunking or diffusion-based approaches [6, 4] with large neural networks, which can lead to unstable learning when combined with conventional RL methods. Several works address these challenges [44, 45]. IBRL [46] trains an imitation policy and then uses it to propose actions for exploration and to bootstrap target values. PA-RL [30] sidesteps applying RL to complex model architectures by learning a Q-function to optimize actions instead of the policy. Other approaches [47, 48, 49] specifically adapt diffusion- or flow-based policies, such as DSRL [50], which runs RL over the latent-noise space of a frozen base diffusion policy and trains a policy to output that noise. Existing methods often couple the algorithm design with the BC policy structure, limiting flexibility across policy classes.
Residual RL and Sample Efficiency. Residual RL [33, 34, 35, 36] provides an alternative approach of learning corrections on top of a fixed base policy, rather than optimizing the entire policy end-to-end. Recent work like ResiP [12] combines a modern action-chunked BC policy with a single-step RL residual policy, although its on-policy RL algorithm has too high sample complexity to be deployed in the real world.
@@
Real-World RL. In the category of real-world RL, QT-Opt[57] used data from a fleet of 7 robots running over the span of weeks to teach a robotic arm with a parallel jaw gripper to grasp diverse objects. More recently, SERL [58] demonstrated insertion and pick-and-place policies by combining a few demonstrations with less than an hour of online interaction. A follow-up work[59] demonstrated more complex behaviors and bimanual control by using a human-in-the-loop approach to overcome the exploration problem. Multiple works have demonstrated RL in the real world for single-arm robots with grippers [60, 46, 50, 61], including [62], who showed that using Q-functions to guide batch online RL is more performant than imitation-based methods. On higher-DOF robots for locomotion [24, 25] and manipulation [13, 63], people often learn in simulation and rely on sim-to-real transfer to run in the real world.
@@ III METHOD
Our method of off-policy RL for residual fine-tuning (ResFiT) of BC policies is illustrated in Fig. 1. It takes as a starting point a base policy, typically trained with BC on an offline dataset of demonstrations using any algorithm or architecture. In our experiments, we choose a base policy with action chunking. Then, we perform online RL using our sample-efficient off-policy recipe to learn residual corrections on top of the base policy. In the following, we outline the details of the full method, including the design choice and implementation details that are crucial to its success.
@@ III-A Base Policy: Behavior Cloning with Action Chunking
Consider an agent that receives observation $o_{t}$ and performs action $a_{t}$ at each timestep $t$. We first collect a dataset $\mathcal{D}_{\text{demos}}$ of demonstrations, e.g., through human teleoperation in the real world, consisting of successful trajectories $\tau=(o_{0},a_{0},o_{1},a_{1},\dots)$. Given this dataset, we train a base policy $\pi_{\psi}(a_{t:t+k}|o_{t})$ with behavior cloning to predict a sequence of $k$ future actions at each timestep. We train the policy to maximize the log-likelihood of the action chunks from the demonstrations, i.e., $\min_{\psi}-\sum_{o_{t},a_{t:t+k}\in\mathcal{D}_{\text{demos}}}\log\pi_{\psi}(a_{t:t+k}|o_{t})$.
Predicting a sequence of actions at each timestep instead of a single action is known to improve performance [64, 5, 4, 65] and alleviate compounding errors in imitation learning by effectively reducing the task horizon of the problem.
Note that although our policies receive only proprioception and image observations $o_{t}$ as input, and do not have access to the true underlying system state $s_{t}$, the remainder of this section will use the state notation for simplicity.
@@ IV-A Simulation Experiments
We evaluate ResFiT’s real-world tractability via simulation experiments using realistic constraints: a single simulation environment, image and robot joint state observations (no privileged object state information), and sparse binary rewards. We test on tasks from Robomimic [74] and DexMimicGen [13], with varying robots, control modes, degrees of freedom, task horizons, and precision requirements. Our experiments use the Robosuite environment [75] built on MuJoCo [76], and we operate at 20 Hz control frequency.
@@
Baselines and Ablations. First, we examine how off-policy residual RL on top of an action-chunked base policy compares to directly performing off-policy RL to learn a single-action policy, starting from the same demos. For this, we chose the state-of-the-art algorithm RLPD [56], an off-policy approach that includes both offline demos and online data in each training batch, and uses layer norm to help alleviate the overestimation of the value function. However, with default settings, it is unable to solve our high-DoF tasks when learning from images. Instead, we create an optimized version of RLPD by incorporating the same off-policy RL design decisions (see Sec. III-C) that we use in our method. We refer to this off-policy RL implementation as “Tuned RLPD”. This means that the same number of demos are used, the same off-policy RL algorithm is used, and the same vision encoders and network architectures are used; the main difference is whether a base policy is being used. For another baseline of using demos but not performing residual learning, we compare to IBRL [46], which uses a pre-trained BC policy during RL to propose actions and bootstrap target values. Finally, we include an online BC fine-tuning baseline called “Filtered BC,” which starts with the same base policy as ResFiT, but does not use residual RL to perform the fine-tuning. It fine-tunes by iteratively performing rollouts and adding successes back into the dataset for continued BC training. This approach is similar in spirit to reward-weighted regression [77] (with 0/1 weights) and self-imitation [78, 79], which prior work has shown can improve performance over the demo-only case [2, 65, 80, 81].
@@ V-A Simulation Results
We evaluate various aspects of ResFiT on simulated benchmark tasks, consisting of both single-arm and bimanual manipulation tasks. On the simulated BoxCleanup task, we first compare our off-policy residual RL recipe to an existing approach [12] of performing residual RL in sim with the on-policy PPO [83] algorithm. Here, with our approach converging at 200k steps versus 40M steps, we see a $\mathord{\sim}$$200\times$ boost in sample efficiency, demonstrating the need for off-policy approaches when considering performing the RL directly in the real world.

### v2

@@ Abstract
Recent advances in behavior cloning (BC) have enabled impressive visuomotor control policies. However, these approaches are limited by the quality of human demonstrations, the manual effort required for data collection, and the diminishing returns from offline data. In comparison, reinforcement learning (RL) trains an agent through autonomous interaction with the environment and has shown remarkable success in various domains. Still, training RL policies directly on real-world robots remains challenging due to sample inefficiency, safety concerns, and the difficulty of learning from sparse rewards for long-horizon tasks, especially for high-degree-of-freedom (DoF) systems.
We present a recipe that combines the benefits of BC and RL through a residual learning framework. Our approach leverages BC policies as black-box bases and learns lightweight per-step residual corrections via sample-efficient off-policy RL. We demonstrate that our method requires only sparse binary reward signals and can effectively improve manipulation policies on high-degree-of-freedom (DoF) systems in both simulation and the real world. In particular, we demonstrate, to the best of our knowledge, the first successful real-world RL training on a humanoid robot with dexterous hands. Our results demonstrate state-of-the-art performance in various vision-based tasks, pointing towards a practical pathway for deploying RL in the real world.
@@ II RELATED WORK
Behavior Cloning. BC has shown success in various settings, from autoregressive next-token prediction [1, 24, 37, 38] and diffusion policies [4, 39, 6, 40, 7], to large transformers that directly output robot actions [5, 41, 42]. Recent advances [43, 8, 9] leverage large transformers, diffusion and flow matching heads, and action chunking to handle long-horizon tasks and achieve impressive performance. Despite these advances, BC has fundamental limitations. The manual effort and infrastructure needed to scale up data collection to these levels are incredibly high, and recent work has shown diminishing returns and performance plateaus with increasing data [6, 13, 12].
Fine-tuning BC Policies with RL. RL fine-tuning has been explored as a method to enhance BC policies by allowing the agent to interact directly with the environment. However, directly fine-tuning modern BC architectures with RL presents significant challenges. State-of-the-art BC policies typically use action-chunking or diffusion-based approaches [6, 4] with large neural networks, which can lead to unstable learning when combined with conventional RL methods. Several works address these challenges [44, 45]. IBRL [46] trains an imitation policy and then uses it to propose actions for exploration and to bootstrap target values. PA-RL [30] sidesteps applying RL to complex model architectures by learning a Q-function to optimize actions instead of the policy. Other approaches [47, 48, 49] specifically adapt diffusion- or flow-based policies, such as DSRL [50], which runs RL over the latent-noise space of a frozen base diffusion policy and trains a policy to output that noise. Existing methods often couple the algorithm design with the BC policy structure, limiting flexibility across policy classes.
Residual RL and Sample Efficiency. Residual RL [33, 34, 35, 36] provides an alternative approach of learning corrections on top of a fixed base policy, rather than optimizing the entire policy end-to-end. Recent work, such as ResiP [12], combines a modern action-chunked BC policy with a single-step RL residual policy; however, its on-policy RL algorithm has too high sample complexity to be deployed in the real world.
@@
Real-World RL. In the category of real-world RL, QT-Opt[57] utilized data from a fleet of 7 robots operating over the course of several weeks to train a robotic arm with a parallel jaw gripper to grasp diverse objects. More recently, SERL [58] demonstrated insertion and pick-and-place policies by combining a few demonstrations with less than an hour of online interaction. A follow-up work[59] demonstrated more complex behaviors and bimanual control by using a human-in-the-loop approach to overcome the exploration problem. Multiple works have demonstrated RL in the real world for single-arm robots with grippers [60, 46, 50, 61], including [62], who showed that using Q-functions to guide batch online RL is more performant than imitation-based methods. On higher-DOF robots for locomotion [24, 25] and manipulation [13, 63], people often learn in simulation and rely on sim-to-real transfer to run in the real world. For manipulation, CASHER [64] trains policies in crowdsourced digital twins, achieving zero-shot transfer with sublinear scaling of human effort. However, it is limited to simple manipulation due to challenges in contact simulation.
@@ III METHOD
Our method of off-policy RL for residual fine-tuning (ResFiT) of BC policies is illustrated in Fig. 1. It takes as a starting point a base policy, typically trained with BC on an offline dataset of demonstrations using any algorithm or architecture. In our experiments, we chose a base policy with action chunking. Then, we perform online RL using our sample-efficient off-policy recipe to learn residual corrections on top of the base policy. In the following sections, we outline the details of the method, including the design choices and implementation details that are crucial to its success.
@@ III-A Base Policy: Behavior Cloning with Action Chunking
Consider an agent that receives observation $o_{t}$ and performs action $a_{t}$ at each timestep $t$. We first collect a dataset $\mathcal{D}_{\text{demos}}$ of demonstrations, e.g., through human teleoperation in the real world, consisting of successful trajectories $\tau=(o_{0},a_{0},o_{1},a_{1},\dots)$. Given this dataset, we train a base policy $\pi_{\psi}(a_{t:t+k}|o_{t})$ using behavior cloning to predict a sequence of $k$ future actions at each timestep. We train the policy to maximize the log-likelihood of the action chunks from the demonstrations, i.e., $\min_{\psi}-\sum_{o_{t},a_{t:t+k}\in\mathcal{D}_{\text{demos}}}\log\pi_{\psi}(a_{t:t+k}|o_{t})$.
Predicting a sequence of actions at each timestep, rather than a single action, is known to improve performance [65, 5, 4, 66] and alleviate compounding errors in imitation learning by effectively reducing the task horizon of the problem.
Note that although our policies receive only proprioception and image observations $o_{t}$ as input, and do not have access to the actual underlying system state $s_{t}$, the remainder of this section will use the state notation for simplicity.
@@ IV-A Simulation Experiments
We evaluate ResFiT’s real-world tractability via simulation experiments using realistic constraints: a single simulation environment, image and robot joint state observations (no privileged object state information), and sparse binary rewards. We test our approach on tasks from Robomimic [75] and DexMimicGen [13], which vary in terms of robots, control modes, degrees of freedom, task horizons, and precision requirements. Our experiments utilize the Robosuite environment [76], which is built on MuJoCo [77], and we operate at a control frequency of 20 Hz.
@@
Baselines and Ablations. First, we examine how off-policy residual RL on top of an action-chunked base policy compares to directly performing off-policy RL to learn a single-action policy, starting from the same demos. For this, we chose the state-of-the-art algorithm RLPD [56], an off-policy approach that includes both offline demos and online data in each training batch, and uses layer norm to help alleviate the overestimation of the value function. However, with default settings, it is unable to solve our high-DoF tasks when learning from images. Instead, we create an optimized version of RLPD by incorporating the same off-policy RL design decisions (see Sec. III-C) that we use in our method. We refer to this off-policy RL implementation as “Tuned RLPD.” This means that the same number of demos are used, the same off-policy RL algorithm is used, and the same vision encoders and network architectures are used; the main difference is whether a base policy is being used. For another baseline of using demos but not performing residual learning, we compare to IBRL [46], which uses a pre-trained BC policy during RL to propose actions and bootstrap target values. Finally, we include an online BC fine-tuning baseline called “Filtered BC,” which starts with the same base policy as ResFiT, but does not use residual RL to perform the fine-tuning. It fine-tunes by iteratively performing rollouts and adding successes back into the dataset for continued BC training. This approach is similar in spirit to reward-weighted regression [78] (with 0/1 weights) and self-imitation [79, 80], which prior work has shown can improve performance over the demo-only case [2, 66, 81, 82].
@@ V-A Simulation Results
We evaluate various aspects of ResFiT on simulated benchmark tasks, consisting of both single-arm and bimanual manipulation tasks. On the simulated BoxCleanup task, we first compare our off-policy residual RL recipe to an existing approach [12] of performing residual RL in sim with the on-policy PPO [84] algorithm. As shown in Fig. 3, our approach converging at 200k steps versus 40M steps, we see a $\mathord{\sim}$$200\times$ boost in sample efficiency, demonstrating the need for off-policy approaches when considering performing the RL directly in the real world.
@@ VI DISCUSSION
ACKNOWLEDGMENTS
We thank Himanshu Gaurav Singh for technical discussions, helping with the filtered BC baseline, and feedback on the paper draft. Younggyo Seo provided valuable input for real-world implementation and offered constructive feedback on the paper. We also thank Carmelo Sferrazza and Ruihan Yang for their comments on the manuscript. Benjamin Colby, Hassan Farooq, and Sumedh Sontakke helped with real-world experiments. Finally, we appreciate the technical discussions during early project development from Quiang Li, Seohong Park, Perry Dong, Hengyuan Hu, and Suvir Mirchandani.